# A modified recursive transfer matrix algorithm for radiation and scattering computation of multilayer spheres

### v1

@@ Abstract
We discusses the electromagnetic scattering and radiation problems of multilayered spheres, reviewing the historical expansion of the Lorentz-Mie theory and the numerical stability issues encountered in handling multilayered spheres. By combining recursive methods with the transfer matrix method, we propose a modified transfer matrix algorithm designed for the stable and efficient calculation of electromagnetic scattering coefficients of multilayered spheres. The new algorithm simplifies the recursive formulas by introducing Debye potentials and logarithmic derivatives, effectively avoiding numerical overflow issues associated with Bessel functions under large complex variables. Numerical test results demonstrate that this algorithm offers superior stability and applicability when dealing with complex cases such as thin shells and strongly absorbing media.
Keywords transfer matrix  $\cdot$ recursive algorithm  $\cdot$ multilayer sphere  $\cdot$ Bessel functions
@@ 1 Introduction
As early as 1951, Aden and Kerker[5] extended the Lorentz-Mie theory to the case of coated spheres. For multilayered spheres, the electric and magnetic fields are represented as multipole expansions, with the coefficients of these expansions determined by the boundary conditions at the interfaces of the sphere’s layers. Kerker[6] further extended this theory to multilayered spheres. In the early days, applying Kerker’s formulas to calculate the scattering of multilayered spheres often encountered various numerical issues, partly due to the use of incorrect recurrence formulas for calculating Bessel functions. Even with correct recurrence formulas, numerical problems still arose when calculating thin-shell spheres, strongly absorbing media, and highly multilayered cases. To address these numerical stability issues, Toon and Ackerman[7] reformulated Kerker’s equations using logarithmic derivatives, providing a stable algorithm for scattering by double-layered spheres. Further, Wu and Wang[8] proposed a concise recursive method to calculate electromagnetic scattering coefficients for arbitrary multilayered spheres. Johnson[9] also developed a slightly different recurrence algorithm. However, numerical errors persisted when dealing with thin, strongly absorbing shells. Kaiser and Schweiger[10] proposed a numerically stable algorithm for calculating Mie scattering coefficients of coated spheres, but it could not be easily extended to arbitrary multilayered spheres. Building on the work of Wu and Wang, Yang[11] proposed an improved recursive algorithm for stable calculations of Mie scattering coefficients in complex cases, such as multilayered and strongly absorbing media. In the field of electromagnetics, Chew[1] developed a layered medium framework suitable for spherical cases, which resulted in concise recurrence formulas. Recently, Yuan et al.[12] introduced these ideas into the spherical layered medium framework, yielding a set of numerically stable calculation schemes. In mathematical physics, Moroz[13] proposed the transfer matrix method for solving the electromagnetic wave scattering problem of multilayered spheres. The transfer matrix method is an important and highly efficient technique for solving general scattering problems, although it still has some limitations in terms of applicability[14, 15].
@@
To expand its applicability, this paper discusses an improved transfer matrix algorithm for the efficient and stable calculation of electromagnetic scattering and radiation by multilayered spheres. Our derivation of the spherical transfer matrix differs slightly from Moroz’s approach, as we use Debye potentials to simplify the expressions, providing a new recursive format and a numerically stable algorithm that avoids overflow issues. Section 2 introduces the transfer matrix method for electromagnetic wave scattering by multilayered spheres and presents new recursive formulas. Section 3 demonstrates a stable calculation scheme for the recursive formulas obtained through logarithmic derivatives or ratios of Bessel functions, with the final algorithm effectively avoiding round-off errors and numerical overflow issues caused by the imaginary part of large complex variables. Section 4 discusses three numerical test cases, showcasing the stability and applicability of the proposed algorithm. Section 5 concludes with a summary of the algorithm and results.
@@ 2 Theory
In spherical coordinates, the complete electromagnetic fields can be expressed using two scalar functions (Debye potential). The Deybe potentials $\pi_{m}$ and $\pi_{e}$ satisfy the scalar Helmholtz equation,
@@
with the Wronskian,
$w_{l}(k_{i}r)=\psi_{l}(k_{i}r)\xi^{\prime}_{l}(k_{i}r)-\xi_{l}(k_{i}r)\psi^{\prime}_{l}(k_{i}r)$ (6)
Furthermore, $\Lambda^{\sigma}_{i}$ is a diagonal matrix, and its specific form depends on the convention. For transverse magnetic (TM) polarization,
@@
The transverse electric (TE) $\Lambda^{e}_{i}$ can be obtained by the duality principle. The matrix $\Lambda^{\sigma}_{i}$ acts as a scaling factor. For the sake of simplicity in the derivation, we can incorporate it into the fundamental matrix, such as $\frac{\epsilon_{i}}{k_{i}}\psi_{l}(k_{i}r_{i})\to\tilde{\psi}_{l}(k_{i}r_{i})$. Without affecting understanding, we omit the superscript $\tilde{\cdot}$ and eventually restore the final expression. Multiplying both sides of the equation by $\Phi^{-1}_{i+1}$, we get
@@
with $z_{i}=k_{i}r_{i}$, $z^{+}_{i}=k_{i+1}r_{i}$. For an $n$-layered sphere, we have the following formula
@@
Since the third kind of Bessel functions are singular at the origin, we set $A_{1}=1$ and $B_{1}=0$. Therefore, the coefficients in the outermost medium are directly given by elements of $M$, i.e., $A_{n+1}=M_{11}$ and $B_{n+1}=M_{21}$. The total scattering coefficients for the $n$-layered sphere are given by the following formula:
$\mathcal{T}_{\sigma,l}=\frac{B_{n+1}}{A_{n+1}}=\frac{M_{11}}{M_{21}}$ (10)
@@
Here, $x=2\pi R/\lambda$ is the size parameter, where $R$ is the radius of the outermost layer of the sphere, and $\lambda$ is the wavelength in vacuum. The absorption efficiency is given by $Q_{abs}=Q_{ext}-Q_{sca}$. The above expression is also applicable to irregularly shaped objects, such as those discussed in our previous work[16] on scattering from randomly shaped geometries. The energy emissivity of the multilayered sphere is given by the following formula[4, 17]:
[…]
Figure 1: Schematic of an n-layer sphere embedded in an infinite dielectric medium. The outer radius, permittivity, and permeability of each layer are $r_{i}{\hskip 0.0pt},\epsilon_{i}{\hskip 0.0pt}$, and $\mu_{i}$, respectively, where $i=1,2,…,n$. The permittivity of the medium is the permittivity in vacuum $\epsilon_{n+1}{\hskip 0.0pt}=\epsilon_{0}$.
@@
Figure 2: The extinction efficiency $Q_{ext}$of the coated sphere as a function of the outermost layer’s size parameter is shown. The plot compares the results with those obtained using RTMA. The refractive indices of the core and shell are $m_{1}=1.33+i0.00$ and $m_{2}=1.33+i$i, respectively.
Figure 3: The extinction efficiency $Q_{ext}$ of the multilayer sphere as a function of the number of layers is shown. The plot compares the results with those obtained using RTMA. The refractive index of each layer is randomly generated.
Figure 4: The emitted energy of a SiC+gold coated sphere at a temperature of 300K, as a function of the outer radius R, is normalized with respect to the Stefan-Boltzmann result. Different curves correspond to different ratios of inner to outer layer radii.
@@ 3 Algorithm
Next, we derive the recursive algorithm. The scattering coefficient for each layer can be given by the following equation:
@@
Substituting the elements of the transfer matrix into the above expression, we have:
@@
When the refractive index of the sphere is real, the arguments of the Bessel function and its derivative are also real, making it easy to obtain robust recurrence formulas for their evaluation. However, when the refractive index is complex, the calculation of the Bessel function becomes unstable as it enters the exponential domain and goes out of bounds. To avoid this ill-conditioning, the calculation formula will involve only the ratios and logarithmic derivatives of the Bessel functions. The logarithmic derivatives are
@@
$\tilde{\mathcal{T}}^{i}_{\sigma,l}=\frac{\psi_{l}(z_{i})}{\xi_{l}(z_{i})}\mathcal{T}^{i}_{\sigma,l}$
@@
$\tilde{\mathcal{T}}^{i}_{\sigma,l}$ $=$ $-\frac{\big{(}F_{l}(z^{+}_{i})-F_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}D_{l}(z^{+}_{i})-D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{\sigma,l}}{\big{(}F_{l}(z^{+}_{i})-D_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}D_{l}(z^{+}_{i})-D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{\sigma,l}}$ (18)
@@
$\tilde{\mathcal{T}}^{i}_{m,l}$ $=$ $-\frac{\big{(}\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{m,l}}{\big{(}\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{m,l}}$
@@
$\tilde{\mathcal{T}}^{i}_{e,l}$ $=$ $-\frac{\big{(}\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{e,l}}{\big{(}\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big{)}+Q^{i+1}_{l}\big{(}\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big{)}\tilde{\mathcal{T}}^{i+1}_{e,l}}$
@@
The main contribution of this paper is the improvement of the transfer matrix method, making it suitable for recursive computation, which can be applied to cases with an ultra-high number of layers and strongly absorbing media. Compared to the work of Wu, Yang, and others, our recursive formulas are more concise and allow for the sequential computation of expansion coefficients for each spherical layer, making the algorithm more efficient. Additionally, we have derived recursive formulas applicable to the outgoing wave scenario.
Bessel functions are frequently used in the analysis of scattering and radiation problems involving electromagnetic waves, acoustic waves, and other phenomena. Numerous studies have been conducted on the precise computation of spherical Bessel functions. Classical algorithms include Miller’s algorithm[19], which utilizes a simple backward recurrence relation, Gautschi’s algorithm[20], and Olver’s algorithm[21], which can automatically estimate truncation errors. These algorithms are capable of computing a class of special functions, including Bessel functions. Some methods[22, 23] for computing Bessel functions are special cases of the aforementioned algorithms. As an auxiliary work, this paper extends a previously proposed stable algorithm [24] for computing Bessel functions to allow the computation of general second-order difference equations. For a general three-term recurrence difference equation, the expression is given by:
$a_{n}\psi_{n-1}+b_{n}\psi_{n}+c_{n}\psi_{n+1}=d_{n}$ (23)
As pointed out by[24], Miller’s algorithm and Gautschi’s algorithm are mathematically equivalent to $LU$ decomposition, while Olver’s algorithm is mathematically equivalent to $UL$ decomposition. We propose a new recursive algorithm based on $LDM^{T}$ decomposition, where $L$, $M^{T}$ and $D$ are lower triangular, upper triangular and diagonal matrices, respectively. Introduce the auxiliary variables, $\gamma_{n}$
$\gamma_{n}$ $=$ $\Big{(}b_{n}-\gamma_{n-1}a_{n}c_{n-1}\Big{)}^{-1}$ (24)
with $\gamma_{0}=0$. The forward and backward substitution algorithms are,
$u_{1}=\gamma_{1}d_{1},\qquad u_{n}=\gamma_{n}(d_{n}-a_{n}u_{n-1})$ (25)
and
$\psi_{\nu}=u_{\nu},\qquad\psi_{n}$ $=$ $u_{n}-\gamma_{n}c_{n}\psi_{n+1}$ (26)
with $\nu$ is the truncation order. The main advantage of this algorithm is that the truncation error could be estimated automatically. To estimate the optimal value of $\nu$, we increase the truncation order from $\nu$ to $\nu+1$. Thus, we obtain the following expression:
$\psi^{(\nu+1)}_{n}-\psi^{(\nu)}_{n}$ $=$ $-\gamma_{n}c_{n}(\psi^{(\nu+1)}_{n+1}-\psi^{(\nu)}_{n+1})$ (27)
$=$ $(-\gamma_{n}c_{n})(-\gamma_{n+1}c_{n+1})\cdots(-\gamma_{\nu-1}c_{\nu-1})(\psi^{(\nu+1)}_{\nu}-\psi^{(\nu)}_{\nu})$
and
$\psi^{(\nu+1)}_{\nu}-\psi^{(\nu)}_{\nu}=u_{\nu}-\gamma_{\nu}c_{\nu}\psi^{(\nu+1)}_{\nu+1}-u_{\nu}=-\gamma_{\nu}c_{\nu}u_{\nu+1}.$ (28)
that is,
$\psi^{(\nu+1)}_{n}-\psi^{(\nu)}_{n}=u_{\nu+1}\prod^{\nu}_{i=n}(-\gamma_{i}c_{i})$ (29)
Let the truncated order be the smallest integer that satisfies the following inequality.
$\Big{|}u_{\nu+1}\prod^{\nu}_{i=n}(-\gamma_{i}c_{i})\Big{|}<\frac{1}{2}\times 10^{-D}$ (30)
The Bessel function $\psi_{n}$ can be computed to a precision of $D$ decimal digits, where $D$ is determined as the truncation criterion for the computational algorithm. The truncation error can be estimated with,
$\epsilon^{(\nu)}=\psi_{n}-\psi^{(\nu)}_{n}=\sum^{\infty}_{s=\nu}u_{s+1}\prod^{s}_{i=n}(-\gamma_{i}c_{i})$ (31)
For the computation of spherical Bessel functions, $a_{n}=c_{n}=1,b_{n}=-(2n+1)/z,d_{n}=0$. In this case, the factorization $LDM^{T}$ simplifies considerably since the underlying matrix is symmetric with $M=L$. This yields the $LDL^{T}$ factorization. With respect to the $LU$ factorization, the computational cost decreases to one half.

### v2

# A modified recursive transfer matrix algorithm for radiation and scattering computation of a multilayered sphere

@@ Abstract
We discuss the electromagnetic scattering and radiation problems of multilayered spheres, reviewing the history of the Lorentz-Mie theory and the numerical stability issues encountered in handling multilayered spheres. By combining recursive methods with the transfer matrix method, we propose a modified transfer matrix algorithm designed for the stable and efficient calculation of electromagnetic scattering coefficients of multilayered spheres. The new algorithm simplifies the recursive formulas by introducing Debye potentials and logarithmic derivatives, effectively avoiding numerical overflow issues associated with Bessel functions under large complex variables. Moreover, by adopting a hybrid recursive strategy, this algorithm can resolve the singularity problem associated with logarithmic derivatives in previous algorithms. Numerical test results demonstrate that this algorithm offers superior stability and applicability when dealing with complex cases such as thin shells and strongly absorbing media.
keywords:
transfer matrix method , recursive algorithm , multilayered sphere , Mie theory
journal: JQSRT
[…]
@@ 1 Introduction
As early as 1951, Aden and Kerker [5] extended the Lorentz-Mie theory to the case of coated spheres. For multilayered spheres, the electric and magnetic fields are represented as multipole expansions, with the coefficients of these expansions determined by the boundary conditions at the interfaces of the sphere’s layers. Kerker [6] further extended this theory to multilayered spheres. In the early days, applying Kerker’s formulas to calculate the scattering of multilayered spheres often encountered various numerical issues, partly due to the use of incorrect recurrence formulas for calculating Bessel functions. Even with correct recurrence formulas, numerical problems still arose when calculating thin-shell spheres, strongly absorbing media, and highly multilayered cases. To address these numerical stability issues, Toon and Ackerman [7] reformulated Kerker’s equations using logarithmic derivatives, providing a stable algorithm for scattering by double-layered spheres. Further, Wu and Wang [8] proposed a concise recursive method to calculate electromagnetic scattering coefficients for arbitrary multilayered spheres. Johnson [9] also developed a slightly different recurrence algorithm. However, numerical errors persisted when dealing with thin, strongly absorbing shells. Kaiser and Schweiger [10] proposed a numerically stable algorithm for calculating Mie scattering coefficients of coated spheres, but it could not be easily extended to arbitrary multilayered spheres. Building on the work of Wu and Wang [8], Yang [11] proposed an improved recursive algorithm for stable calculations of Mie scattering coefficients in complex cases, such as multilayered and strongly absorbing media. In the field of electromagnetics, Chew [1] developed a layered medium framework suitable for spherical cases, which resulted in concise recurrence formulas. Recently, Yuan et al. [12] introduced these ideas into the spherical layered medium framework, yielding a set of numerically stable calculation schemes. In mathematical physics, Moroz [13] introduced the transfer matrix method to address the electromagnetic wave scattering problem in multilayered spheres. This approach, referred to as the Recursive Transfer Matrix Method (RTMM) by Moroz, is denoted here as the Recursive Transfer Matrix Algorithm (RTMA) to emphasize its algorithmic implementation. The transfer matrix method is an important and highly efficient technique for solving general scattering problems, although it still has some limitations in terms of applicability [14, 15].
@@
To expand its applicability, this paper discusses an improved transfer matrix algorithm for the efficient and stable calculation of electromagnetic scattering and radiation by multilayered spheres. Our derivation of the spherical transfer matrix differs slightly from Moroz’s approach, as we use Debye potentials to simplify the expressions, providing a new recursive format and a numerically stable algorithm that avoids overflow issues. Section 2 introduces the transfer matrix method for electromagnetic wave scattering by a multilayered sphere and presents new recursive formulas. Section 3 demonstrates a stable calculation scheme for the recursive formulas obtained through logarithmic derivatives or ratios of Bessel functions, with the final algorithm effectively avoiding round-off errors and numerical overflow issues caused by the imaginary part of large complex variables. Section 4 discusses three numerical test cases, showcasing the stability and applicability of the proposed algorithm. Section 5 concludes with a summary of the algorithm and results.
@@ 2 Theory
In spherical coordinates, the electromagnetic fields can be decomposed into transverse electric (TE) waves and transverse magnetic (TM) waves. For TE waves, $E^{\rm TE}=\nabla\times(\mathbf{r}\,\pi_{m}).$ For TM waves, $H^{\rm TM}=\nabla\times(\mathbf{r}\,\pi_{e}).$ The two scalar functions $\pi_{m}$ and $\pi_{e}$ are the so-called Debye potentials, which satisfy the scalar Helmholtz equation.
@@
with the Wronskian of $\psi_{l}(k_{i}r)$ and $\xi_{l}(k_{i}r)$,
$w_{l}(k_{i}r)=\psi_{l}(k_{i}r)\xi^{\prime}_{l}(k_{i}r)-\xi_{l}(k_{i}r)\psi^{\prime}_{l}(k_{i}r)=\mathrm{i}$ (6)
with $\mathrm{i}=\sqrt{-1}$. Furthermore, $\Lambda^{\sigma}_{i}$ is a diagonal matrix, and its specific form depends on the convention. For transverse magnetic (TM) polarization,
@@
The transverse electric (TE) $\Lambda^{e}_{i}$ can be obtained by the duality principle.
The matrix $\Lambda^{\sigma}_{i}$ acts as a scaling factor. For the sake of simplicity in the derivation, we can incorporate it into the fundamental matrix, such as $\frac{\epsilon_{i}}{k_{i}}\psi_{l}(k_{i}r_{i})\to\tilde{\psi}_{l}(k_{i}r_{i})$. Without affecting understanding, we omit the superscript $\tilde{\cdot}$ and eventually restore the final expression. Multiplying both sides of the equation by $\Phi^{-1}_{i+1}$, we get
@@
[…]
For an $n$-layer sphere, we have the following formula
@@
Since the third kind of Bessel functions are singular at the origin, we set $A_{1}=1$ and $B_{1}=0$. Therefore, the coefficients in the outermost medium are directly given by the matrix elements of $2\times 2$ matrix $M$, i.e., $A_{n+1}=M^{11}$ and $B_{n+1}=M^{21}$. The total scattering coefficients for the $n$-layer sphere are given by the following formula:
$\mathcal{T}_{\sigma,l}=\frac{B_{n+1}}{A_{n+1}}=\frac{M^{11}}{M^{21}}$ (10)
@@
Here, $x=2\pi R/\lambda$ is the size parameter, $R$ is the radius of the outermost layer of the sphere, and $\lambda$ is the wavelength in vacuum. The absorption efficiency is given by $Q_{abs}=Q_{ext}-Q_{sca}$. The above expression is also applicable to irregularly shaped objects, such as those discussed in our previous work [16] on scattering from randomly shaped geometries, and spheres embedded in an absorbing medium [17]. The energy emissivity of the multilayered sphere is given by the following formula [4, 18]:
$H_{s}$ $=$ $\frac{8\pi^{2}R^{2}hc^{2}}{\lambda^{5}}\int^{\infty}_{0}d\lambda\frac{1}{\exp{\frac{hc}{k_{B}T\lambda}}-1}Q_{abs}$ (13)
(a) Standing wave scenario
(b) Outgoing wave scenario
Figure 1: Schematic of an $n$-layer sphere embedded in an infinite dielectric medium. The outer radius, permittivity, and permeability of each layer are $r_{i}$, $\epsilon_{i}$, and $\mu_{i}$, respectively, with i=1,2,,ni=1,2,...,n. The surrounding medium has vacuum permittivity $\epsilon_{n+1}=\epsilon_{0}$.
@@ 3 Algorithm
In this section, we present a recursive algorithm based on the transfer matrix method, referred to as the modified Recursive Transfer Matrix Algorithm (mRTMA). It is important to note that while the RTMA involves the combined use of the transfer matrix, it does not align with the definition of a recursive algorithm as commonly understood in the field of scientific computing.
The scattering coefficient for each layer can be given by the following equation:
@@
with
$\mathcal{T}^{0}_{\sigma,l}=0.$
Substituting the elements of the transfer matrix (eq.8) into the above expression, we have:
@@
When the refractive index of the sphere is real, the arguments of the Bessel function and its derivative are also real, making it easy to obtain robust recurrence formulas for their evaluation. However, when the refractive index is complex, the calculation of the Bessel function becomes unstable as it enters the exponential domain and goes out of bounds. To avoid this ill-conditioning, the calculation formula will involve only the ratios and logarithmic derivatives of the Bessel functions.
The logarithmic derivatives are
@@
$\breve{\mathcal{T}}^{i}_{\sigma,l}=\frac{\psi_{l}(z_{i})}{\xi_{l}(z_{i})}\mathcal{T}^{i}_{\sigma,l}$
@@
$\breve{\mathcal{T}}^{i}_{\sigma,l}$ $=$ $-\frac{\big(F_{l}(z^{+}_{i})-F_{l}(z_{i})\big)+Q^{i+1}_{l}\big(D_{l}(z^{+}_{i})-D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{\sigma,l}}{\big(F_{l}(z^{+}_{i})-D_{l}(z_{i})\big)+Q^{i+1}_{l}\big(D_{l}(z^{+}_{i})-D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{\sigma,l}}$ (18)
@@
$\breve{\mathcal{T}}^{i}_{m,l}$ $=$ $-\frac{\big(\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z_{i})\big)+Q^{i+1}_{l}\big(\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{m,l}}{\big(\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big)+Q^{i+1}_{l}\big(\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{m,l}}$
@@
$\breve{\mathcal{T}}^{i}_{e,l}$ $=$ $-\frac{\big(\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}F_{l}(z_{i})\big)+Q^{i+1}_{l}\big(\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{e,l}}{\big(\sqrt{\epsilon_{i+1}\mu_{i}}F_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big)+Q^{i+1}_{l}\big(\sqrt{\epsilon_{i+1}\mu_{i}}D_{l}(z^{+}_{i})-\sqrt{\epsilon_{i}\mu_{i+1}}D_{l}(z_{i})\big)\breve{\mathcal{T}}^{i+1}_{e,l}}$
@@
The main contribution of this paper is the improvement of the transfer matrix method, making it suitable for recursive computation, which can be applied to cases with an ultra-high number of layers and strongly absorbing media. Compared to the works of Wu [8] and Yang [11], our recursive formulas are more concise and allow for the sequential computation of expansion coefficients for each spherical layer, making the algorithm more efficient. Additionally, we have derived recursive formulas applicable to the outgoing wave scenario.
We next address potential issues that may arise during the implementation of our algorithm and propose corresponding solutions. For real arguments $z$ (layers with purely real material constants), $\psi_{l}$ is known to exhibit real zeros, leading to singularities in $D_{l}$ as noted by one of the reviewers. To handle this, a hybrid recursive scheme combining Eq. 15 and Eq. 17 can be employed. Specifically, for spherical layer that encounter the singularity of $D_{l}$, Eq. 15 can be applied for computation. This approach represents another advantage of the mRTMA. For complex arguments $z$, $\xi_{l}$ diverges to infinity while $\psi_{l}$ approaches zero as $l$ goes to infinity. However, practical computations involve a truncation order, $l_{max}$, approximately $|z|+p|z|^{1/3}+1$, where $p$ is set to be about 4. Within the truncation range ($l<l_{max}$), the ratio $|\xi_{l}/\psi_{l}|$ increases monotonically. Thus, ensuring the boundedness of $|\xi_{l_{max}}/\psi_{l_{max}}|$ suffices for stability. Empirically, it can be observed that $\xi_{l}$ and $\psi_{l}$ are proportional to $e^{-\text{Im}z}$ and $e^{\text{Im}z}$, respectively, causing $|\xi_{l}/\psi_{l}|$ decay exponentially with $2\text{Im}z$. At the truncation order, $l_{max}$, it is straightforward to verify that $|\xi_{l}/\psi_{l}|<1$.